OpenStudy (lina17):

solve for x: -3(x+3)=-3(x+1)-5 help step by step

OpenStudy (kkutie7):

notice that there isn't anything you can multiply or divide to the other side to make this simpler. you can start by multiplying what is in the parenthesis by the -3 out side of it before continuing.

OpenStudy (lina17):

so it would be -3x+9=-3(x+1)= -5

OpenStudy (serenity74):

-3x-9 = -3x-3-5 -3x-9 = -3x -8 0 = 1

OpenStudy (lina17):

how is it 1 when its -3x+9= -3x-3- -5

OpenStudy (serenity74):

i distributed it and then solved for it

OpenStudy (lina17):

can you go step by step please

OpenStudy (misty1212):

Hi!!

OpenStudy (lina17):

hi

OpenStudy (kkutie7):

now you need to add all common terms can you do that?

OpenStudy (misty1212):

not to butt in but i think @Kkutie7 copied wrong

OpenStudy (lina17):

ya its kind of confusing

OpenStudy (kkutie7):

yes! let me fix that

OpenStudy (misty1212):

this is the quesion right \[-3(x+3)=-3(x+1)-5 \]

OpenStudy (lina17):

yes

OpenStudy (misty1212):

ok i will go now

OpenStudy (kkutie7):

New step one: \[-3(x+3)=-3(x+1)-5\] \[-3x-9=-3x-3-5\] now you can add like terms

OpenStudy (lina17):

so you add -3x+-3x ?

OpenStudy (kkutie7):

first you need to move everything with like term to the same side. \[-3x+3x=-3-5+9\] do you see?

OpenStudy (serenity74):

I did go step by step